OpenStudy (anonymous):

The equation shows one mole of ethanol fuel being burned in oxygen. Convert the energy released into its equivalent mass. C2H5OH(l) + 3 O2(g) Æ 2 CO2(g) + 3 H2O (l) ΔH = -1418 kJ/mol

OpenStudy (anonymous):

E=mc^2

OpenStudy (anonymous):

i still dont undererstand it though...can you do it step by step?

sam (.sam.):

E=energy m=mass c=speed of light

OpenStudy (callisto):

For burning one mole of ethanol, energy released = 1418 kJ By Einstein's mass - energy relationship E=mc^2 1418x10^3 = m (3x10^8)^2 m= (1418x10^3) / (3x10^8)^2=? I hope I'm not incorrect... I haven't done anything like that for a month :(

OpenStudy (anonymous):

E= ΔH = -1418 kJ/mol